A market consists of three consumers whose demand curves are: P= 35-0.5Qa; P= 50-0.25Qb; and P= 40-2Qc. Calculate the market demand for the commodity and the equilibrium output and price if the supply function is given by Qs= 40+3.5P. (8 marks)
1
Expert's answer
2020-08-14T16:48:54-0400

"Qa=\frac{P-35}{-0.5}"

"Qb=\frac{P-50}{-0.25}"

"Qc=\frac{P-40}{-2}"

"Qs=Qa+Qb+Qc"

"\frac{P-35}{-0.5}+\frac{P-50}{-0.25}+\frac{P-40}{-2}=40+3.5P"

"4(P-35)+8(P-50)+P-40=-80-7P"

"20P=500"

"P=25"

"Qs=87.5"